OpenStudy (anonymous):

Find the center and radius of the circle with the following equation 8xsquare + 8y square + 64x-32y+32=0

OpenStudy (anonymous):

\[8x^2+8y^2+64x-32y+32=0\] divide by 8 as step one and get \[x^2+8x+y^2-4y+4=0\] subtract 4 and get \[x^2+8x+y^2-4y=-4\] then "complete the square" to make it look like \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

you know how to complete the square? that is the only hardish part

OpenStudy (anonymous):

(-4,2) 4